OpenStudy (anonymous):

Find the work done in moving an object along a vector r = 3i + 2j - 5k if applied force is F = 2I - J - K

OpenStudy (turingtest):

\[W=\int\vec F\cdot d\vec s\]so it's just the integral of the dot product

OpenStudy (turingtest):

\[W=\int\vec F\cdot d\vec r\]

OpenStudy (turingtest):

but there are no bounds on your integral, so perhaps they just want the dot product...?

OpenStudy (anonymous):

yeah... the question does not have bounds it is just plane like that

OpenStudy (turingtest):

maybe you should post this in the physics group (though I usually would help you in that dept. too) I'm just not sure about the meaning of the question since there is no actual path, just a direction

OpenStudy (anonymous):

ohkay i will do that now

OpenStudy (turingtest):

I'm starting to think they just want the dot product

OpenStudy (turingtest):

because r has a length, so it can be interpreted as a line segment so just do\[W=\vec F\cdot\vec r\]

OpenStudy (turingtest):

...at least that's what I think

OpenStudy (anonymous):

ohkay... because the 1st equation you gave out is for the conservative force field

OpenStudy (turingtest):

?

OpenStudy (turingtest):

the first equation I gave is for the work done by a force function acting on a parametric displacement function

OpenStudy (turingtest):

the force does not need to be conservative for that equation to hold

OpenStudy (turingtest):

it's actually a line integral\[W=\int\limits_C\vec F\cdot d\vec s\]so it doesn't have to be conservative

OpenStudy (turingtest):

if\[\oint\vec F\cdot d\vec s=0\]then \(\vec F\) is conservative

OpenStudy (anonymous):

yes i agree

OpenStudy (turingtest):

but like I say, I don't even think your question wants you to do an integral... as far as I can see

OpenStudy (anonymous):

yep cause it does not have bounds

OpenStudy (maheshmeghwal9):

\[|\overrightarrow F|=\sqrt{(2)^2+(-1)^2+(-1)^2}=\sqrt{6}.\]\[|\overrightarrow r|=\sqrt{(3)^2+(2)^2+(-5)^2}=\sqrt{38}.\]\[W=F*r*\cos \theta\]But I don't know what is theta?

OpenStudy (anonymous):

so where will i get the angle from

OpenStudy (maheshmeghwal9):

I think Dot product of vector will come into application. Only i think.

OpenStudy (turingtest):

you can just do it with the dot product method

OpenStudy (turingtest):

\[\vec v=\langle a,b,c\rangle\]\[\vec u=\langle x,y,z\rangle\]\[\vec u\cdot\vec v=ax+by+cz\]

OpenStudy (turingtest):

that is the definition of the euclidean inner product in \(\mathbb R^3\)

OpenStudy (anonymous):

ohk i will try with the dot product

OpenStudy (turingtest):

another definition*

OpenStudy (maheshmeghwal9):

@TuringTest I wanna ask a question. Plz help:)

OpenStudy (turingtest):

I can try, but please post separately I will help if I know how of course :)

OpenStudy (maheshmeghwal9):

which is true? 1.) \[W=\overrightarrow F* \overrightarrow S* \cos \theta\] 2.) \[W=F*S* \cos \theta.\]

OpenStudy (turingtest):

not sure what the second means exactly, but if F and S are magnitudes then 2) is true

OpenStudy (anonymous):

@TuringTest must derivate the vector r ?

OpenStudy (turingtest):

no, it is a line integral

OpenStudy (turingtest):

...normally I mean

OpenStudy (turingtest):

the formula is force times distance equals work basically

OpenStudy (turingtest):

since r is a distance no derivative required

OpenStudy (anonymous):

so i use the normak r given out ?

OpenStudy (maheshmeghwal9):

Isn't it like this "Work is dot product of force & displacement"? \[W=\overrightarrow F . \overrightarrow S\]\[\implies W= |\overrightarrow F| \times |\overrightarrow S| \times \cos \theta= F*S* \cos \theta\] ???????????????????/

OpenStudy (maheshmeghwal9):

@TuringTest Plz:)

OpenStudy (turingtest):

if it is a straight line, yes but if the path curves you have to parameterize it in terms of time and that method will not work\[\vec u\cdot\vec v=\|\vec u\|\|\vec v\|\cos\theta_{uv}\]

OpenStudy (anonymous):

w = f . r =9 am i right?

OpenStudy (turingtest):

that's the formal definition statement^

OpenStudy (maheshmeghwal9):

oh i see? @kay-jay what value did u take as theta?

OpenStudy (maheshmeghwal9):

theta?

OpenStudy (turingtest):

@kay-jay I get 9 as well

OpenStudy (turingtest):

you do not need theta @maheshmeghwal9 there are two ways to define our dot product

OpenStudy (anonymous):

am not using the dot product that needs an angle

OpenStudy (turingtest):

the formula\[\vec v\cdot\vec u=\|\vec v\|\|\vec u\|\cos\theta\] is more limited than the formula I prescribed above; it only works when the vectors are constant and do not change with time in this case we could use this formula as well, but I forget how to find the angle... gotta look at my notes

OpenStudy (anonymous):

@TuringTest so you think our answer might be true ?

OpenStudy (maheshmeghwal9):

Oh i see what did u both do? That was an extraordinary shot^ Thanx for nice solution @Turingtest :)

OpenStudy (maheshmeghwal9):

It must be true:)

OpenStudy (turingtest):

our answer is true if I have interpreted the problem correctly I am saying that the problem states that the given force F is acting along a vector r, and that the magnitude of vector r represents distance that it travels if I am right about that interpretation I think that we have the right answer

OpenStudy (maheshmeghwal9):

yeah Answer depends on Interpretation.

OpenStudy (turingtest):

PS: the only way I know to find the angle between them is to use the other formula for the dot product\[\vec u\cdot\vec v=ax+by+cz\]so it would be redundant to try to find the angle

OpenStudy (anonymous):

am writing on monday at 14h00 so i want to represent this group with good marks. @TuringTest you are really helping... thank you and thank you also @maheshmeghwal9

OpenStudy (maheshmeghwal9):

yw:D Best of luck for ur representation^_^ @kay-jay

OpenStudy (turingtest):

you could say\[\vec u\cdot\vec v=\|\vec u\|\|\vec v\|\cos\theta\implies\theta=\cos^{-1}{\vec u\cdot\vec v\over\|\vec u\|\|\vec v\|}\]but see how you still have to use\[\vec u\cdot\vec v=ax+by+cz\]to find the angle so like I say, that method is redundant at best

OpenStudy (anonymous):

@TuringTest the one you mension now is for three dimensional

OpenStudy (turingtest):

yes

OpenStudy (maheshmeghwal9):

Wow seems more interesting:D

OpenStudy (anonymous):

mathematics is interesting but challenging

OpenStudy (maheshmeghwal9):

^_^

OpenStudy (turingtest):

I think one attribute contributes to the other :)

OpenStudy (maheshmeghwal9):

yeah i think too:D